Tohban Report 2014-05-21
From RHESSI Wiki
|Start Date:||14 May 2014|
|End Date:||21 May 2014|
|Tohban:||Lindsay Glesener (Vice Tohban: Säm / Milo )|
|Next Tohban:||Juan Carlos (vice tohban Martin)|
|List all reports|
It was a very quiet week, with all B and C flares and no M class flares. The GOES baseline level is mid-B class. There are 11 ARs on the disk. Things are becoming slightly more active (within the last 24 hours), but still only small C flares. STEREO A and B (which give almost the same view now) show that the far side of the Sun is also quiet.
How many GOES flares occurred?
Flares above B, C, M, X class were 3 26 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 2 18 0 0
The SSR fill reached a max of 31% and was at 7% at the end of yesterday's pass set. In other words it didn't quite empty.
CP1 and CP2 are at 118.4K and 116.8K, respectively. These are up a bit but are trending normally with the day cycle.
No data gaps!
Due to attenuator motions last week, the set of detectors used for the livetime calculation was changed by doubling up on D2 (and removing D7). We now are using G1, G2, G2, and G5. The commanding was done May 16 at 2014-136-02:38 UTC. The thin attenuator came back out shortly after. No spurious attenuator motions were noticed since then.
On May 17 at 2014-137-04:03:25 UTC, many thresholds were changed:
- G3 front fast threshold from 0x22 to 0x30
- G6 front fast threshold from 0xF0 to 0xFF (max)
- G1 rear fast threshold from 0x45 to 0x50 (Due to a temporary snafu, this threshold was incorrect for approximately 3 minutes.)
- G4 rear fast threshold from 0x80 to 0x90
G8 appears to have fallen into the same state as G9 with respect to the rear segment, i.e. the rear segment is essentially not functional for data collection. No action was taken on this in the past week, but see the note at the end of this section.
On May 19, an "experiment" was performed on D6, whereby the HV was dropped to 0, stayed off for an orbit, and then brought back up to ~2100V (the same as the starting value) over the next two BGS passes. Some interesting things were found. The detector appeared to desegment at just under 900V on the way down, but segmented at 20-40V on the way back up. (Though Albert notes that this is apparent segmentation only, and doesn't consider the relaxed state after the leakage current has been allowed to settle. In other words the real segmentation / desegmentation voltage might not be exactly these values.) The health of the detector appeared to be better than the health before the "experiment," in the sense that the livetime finished at ~80% (as opposed to almost zero beforehand), though it has slowly declined a bit since then. In this respect, the experiment is considered a success! A writeup of the experiment can be found here.
The exact commands:
- 14-139-17:28:53 /IHVDAC DETECTOR=6, VOLTAGE=88 ; was 108
- 14-139-18:59:13 /IHVDAC DETECTOR=6, VOLTAGE=48 ; was 88
- 14-139-19:00:37 /IHVDAC DETECTOR=6, VOLTAGE=0 ; was 48
- 14-139-20:36:36 /IHVDAC DETECTOR=6, VOLTAGE=10 ; was 0
- 14-139-20:38:27 /IHVDAC DETECTOR=6, VOLTAGE=28 ; was 10
- 14-139-20:40:41 /IHVDAC DETECTOR=6, VOLTAGE=48 ; was 28
- 14-139-22:17:26 /IHVDAC DETECTOR=6, VOLTAGE=68 ; was 48
- 14-139-22:20:14 /IHVDAC DETECTOR=6, VOLTAGE=78 ; was 68
- 14-139-22:22:55 /IHVDAC DETECTOR=6, VOLTAGE=88 ; was 78
- 14-139-23:58:52 /IHVDAC DETECTOR=6, VOLTAGE=98 ; was 88
- 14-140-00:01:21 /IHVDAC DETECTOR=6, VOLTAGE=108 ; was 98
Emboldened by this success, we decided to try the same thing with Detector 9. This test is scheduled for Tuesday. It would be good to try it for D8 as well, but the fact that we can't get real-time events from D8 during the pass due to antenna interference makes D9 the better candidate.
The spin rate is 14.7 rpm and Ops will perform a spinup later today. UPDATE: This was performed today (May 21). We spun up from 14.7 RPM to just shy of 15.0 RPM. The spin up started 2014-141-20:22 UTC and returned to normal mode at 21:32 UTC.
In previous Ops meetings it was noticed that there is some "spiky" data that had come through in the last few weeks. Investigation of this by Jim, Richard, and Brian found that a software fix was necessary in order to properly deal with this data as dropouts instead of spikes. The issue has been corrected and the software fix is described here, in a May 15, 2014 entry.
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 minutes|
|Night time data (rears)||+/- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Detector problems?||Many adjustments to Det settings; see above.|